# On Computational Limits and Provably Efficient Criteria of Visual Autoregressive Models: A Fine-Grained Complexity Analysis

### v1

@@ Appendix D Error Analysis of VQVAE Decoder
Lemma D.1 (Error analysis of Convolution Layer).
@@
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
@@
$\|\phi_{\rm conv}(X)-\phi_{\rm conv}(X^{\prime})\|_{\infty}\leq C\epsilon M$
@@
$\leq$ $~{}\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}\epsilon\cdot M$
$\leq$ $~{}9\cdot c_{\rm in}\epsilon M$
$=$ $~{}C\epsilon M$
@@
Lemma D.2 (Error analysis of Fast VQ-VAE Decoder ).
@@
By Lemma C.7, we have shown that fast attention computation will introduce an approximation error no more than $1/\operatorname{poly}(n)$.
Then, by Lemma C.5 and Lemma D.1, we still can bounded have shown the approximation error by $1/\operatorname{poly}(n)$ after passing another up-interpolation layer, convolutions layer.
@@
Appendix E Running Time
@@
In this section, we conduct the running time analysis of every component of the $\operatorname{\mathsf{VAR}}$ model and the fast $\operatorname{\mathsf{VAR}}$ model. In Section E.1, we conduct the running time analysis of the $\operatorname{\mathsf{VAR}}$ transformer and the fast transformer. In Section E.2, we conduct the running time analysis of the feature map reconstruction block. In Section E.3, we conduct the running time analysis of the VQVAE Decoder and fast VQVAE Decoder.
E.1 Phase 1: Running Time of Token Maps Generation
In this section, we present a lemma on the time complexity of the token maps generation phase of $\operatorname{\mathsf{VAR}}$ and reference the results in [5] to derive a faster variant of $\operatorname{\mathsf{VAR}}$.
@@ E.1 Phase 1: Running Time of Token Maps Generation
Lemma E.1 (Running time of $\operatorname{\mathsf{VAR}}$ Transformer).
@@
  • Let $K$ denote the total number of the token maps.
@@
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
@@
then the time complexity of $\operatorname{\mathsf{VAR}}$ generation is $O(n^{4}d)$.
@@
Firstly, we consider the $k$-th autoregressive token map generation. We use $L_{k}$ to denote the total number of the tokens input into the $\operatorname{\mathsf{VAR}}$ transformer at $k$-th step and $L_{k}$ can be calculated as the following:
$L_{k}=$ $~{}\sum_{i=1}^{k}(\alpha^{i-1})^{2}$
$=$ $~{}\frac{\alpha^{2k}-1}{\alpha^{2}-1}$
$\leq$ $~{}\frac{\alpha^{2k}}{\alpha^{2}-1}$
$\leq$ $~{}\frac{\alpha^{2k}}{0.5\alpha^{2}}$
$=$ $~{}2\cdot\alpha^{2k-2}$ (5)
Thus, at the $k$-th step, we use $X_{k}\in\mathbb{R}^{L_{k}\times d}$ to denote the input matrix. The transformer computation cost at the $k$-th step is $O(L_{k}^{2}d)$. We then sum up the computation time across all $K$ steps:
${\cal T}=$ $~{}O((L_{1}^{2}+L_{2}^{2}+\dots+L_{K}^{2})\cdot d)$
$\leq$ $~{}O(\sum_{k=1}^{\log_{a}{n}+1}(2\cdot\alpha^{2k-2})^{2}\cdot d)$
$=$ $~{}O(\sum_{k=1}^{\log_{a}{n}+1}4\alpha^{4k-4}\cdot d)$
$=$ $~{}O(n^{4}d)$
In the first step, the total time is obtained by summing the times for each $k$-th step, while the second step is a consequence of Eq. (E.1), and the final step is attributed to elementary algebra.
Lemma E.2 (Fast Attention Computation, Theorem 1.4 of [5]).
Let $\mathsf{AAttC}$ be defined as Definition 4.2. Then we have $\mathsf{AAttC}(n,d=\log n,R=(\sqrt{\log n}),\delta=1/\operatorname{poly}(n))$ can be solved in time $\mathcal{T}_{\rm mat}(n,n^{o(1)},d)=n^{1+o(1)}$.
Now we can apply the result Lemma E.2 to the $\operatorname{\mathsf{VAR}}$ Transformer.
Lemma E.3 (Running time of Fast $\operatorname{\mathsf{VAR}}$ Transformer, formal version of Lemma 5.1).
Assuming the conditions below are satisfied:
  • Denote $K$ as the total number of generated token maps.
  • Let $k\in[K]$.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $d=O(\log(n))$ denote embedding size of each token.
Then, the total running time of the $\operatorname{\mathsf{VAR}}$ Transformer for generating token maps can be accelerated to $O(n^{2+o(1)}d)$.
Proof.
To generate the $k$-th token map, let the input of this step be $X\in\mathbb{R}^{L_{k}\times d}$. And we have
$L_{k}\leq 2\cdot\alpha^{2k-2}$ (6)
where this step is a consequence of Eq. (E.1). So the transformer computation cost at the $k$-th step can be improved from $O(L_{k}^{2}d)$ to $O(L_{k}^{1}d)$ by using the result of Lemma E.2. Then, we sum up the computation time across all $K$ steps:
${\cal T}=$ $~{}O((L_{1}+L_{2}+\dots+L_{K})\cdot d)$
$=$ $~{}O(n^{2+o(1)}d)$
The computation time in the first step is determined by summing all $k$-th step times, while the second step is a consequence of Eq. (6), and the last step is attributed to elementary algebra. ∎
E.2 Phase 2: Running Time of Feature Map Reconstruction
In this section, we analyze the running time of the feature map reconstruction layer.
Lemma E.4 (Running time of Feature Map Reconstruction Layer, formal version of Lemma 5.2).
Assuming the conditions below are satisfied:
  • Define $K$ as the total number of the token maps.
  • Let $k\in[K]$.
  • Let $d$ denote the embedding size of each token.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
then the total runtime of the $\operatorname{\mathsf{VAR}}$ models for feature map reconstruction is $O(n^{2}d^{2}\log n)$.
Proof.
For each $k\in[K]$, $\operatorname{\mathsf{VAR}}$ Model will up-interpolate token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$ to $r^{\prime}_{k}\in\mathbb{R}^{n\times n\times d}$ by using bicubic interpolation defined in Definition 3.2. Specifically, the computation of each entry of $r^{\prime}_{k}$ requires $32$ multiplications and $16$ additions (see more details in Definition 3.2). Thus, the computation cost for the up-interpolation of each token map is
${\cal T}^{k}_{\rm up}=$ $~{}48\cdot n^{2}d$
@@ E.2 Phase 2: Running Time of Feature Map Reconstruction
There are total $O(\log n)$ token maps needed to be up-interpolated, so the total time for up-interpolation is
@@
The computation time in the first step is determined by summing the $\log(n)$ up-interpolation time, and the second step is a consequence of elementary algebra.
Furthermore, to address the information loss in the up-interpolation process, the $\operatorname{\mathsf{VAR}}$ Model uses a convolution operation $\phi(\cdot)$ on the token map $\{r^{\prime}_{1},\dots,r^{\prime}_{K}\}$ generated by up-interpolation. We assume the convolution kernel size is $3\times 3\times d$, and the convolution layer does not change the dimension of each token map, i.e., for each $i\in[K]$, $\phi(r^{\prime}_{i})\in\mathbb{R}^{n\times n\times d}$. Hence, for every entry in $\phi(r^{\prime}_{i})$, it needs $O(d)$ operations. Then, we can have the convolution computation time for one token map is
@@
In the first step, the total time is obtained by $O(\log(n))$ convolution operations, while the second step results from simple algebra.
@@
The first step is derived by summing the times for up-interpolation operations, convolution operations, and token map additions, while the second step is due to basic algebra. ∎
E.3 Phase 3: Running Time of VQ-VAE Decoder
@@
Lemma E.5 (Running time of VQ-VAE Decoder).
Assuming the conditions below are satisfied:
@@
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
then the total running time of the VQ-VAE Decoder is $O(n^{4}d)$.
@@
Part 1. Running time of Up-interpolation layers. For each $l\in[k_{1}]$, we assume $M^{l}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the output feature map from the $l$-th up-interpolation layer. For every entry of $M^{l}$, it needs $32$ multiplications and $16$ additions (see more details in Definition 3.2). Thus, the computation cost for the feature map $M^{l}$ is
${\cal T}^{l}_{\rm up}=$ $~{}48\cdot O(n)\cdot O(n)\cdot O(d)$
@@
The first step is derived by summing the computation costs for each entry of $M^{l}$, while the second step is due to basic algebra.
@@
${\cal T}^{l}_{\rm attn}=O(n^{4}d)$
@@
${\cal T}_{\rm attn}=$ $~{}k_{2}\cdot O(n^{4}d)$
$=$ $~{}O(n^{4}d)$
The first step is derived by summing the computation costs for each up-interpolation layer, while the second step is due to the fact that $k_{2}$ is a constant.
@@
The first step is derived by summing the computation costs for each entry of $M^{l}$, while the second step stems from basic algebra.
@@
$=$ $~{}O(n^{4}d)$
The first step results from summing the computation costs of up-interpolation, attention, and convolution layers, while the second step is a consequence of $n\gg d$.
@@
Lemma E.6 (Running time of Fast VQ-VAE Decoder, formal version of Lemma 5.3).
In the case that the following conditions are satisfied:
  • Assume $k_{1},k_{2},k_{3}\in\mathbb{N}$ are constant numbers.
  • We represent the input feature map by $X\in\mathbb{R}^{n\times n\times d}$.
  • We assume there are $k_{1}$ up-interpolation layers $\phi_{\rm up}$ defined in Definition 3.2.
@@
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
then the total runtime of the VQ-VAE decoder can be accelerated to $O(n^{2+o(1)}d)$.
@@
As the same in Eq. (E.5) and Eq. (E.5), the computation cost for up-interpolation layers and convolution layers in VQ-VAE decoder still needs $O(n^{2}d)$ and $O(n^{2}d^{2})$, respectively.
For each $l\in[k_{2}]$, we assume $M^{l-1}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the input feature map for the $l$-th attention layer. We can consider the input of the attention layer as a sequence with length $O(n^{2})$ and embedding dimension $O(d)$. By using the result of Lemma E.2, the computation cost of $M^{l}$ can be speed up to
${\cal T}^{l}_{\rm attn}=O(n^{2+o(1)}d)$
@@
$=$ $~{}O(n^{2+o(1)}d)$
The computation cost in the first step is obtained by adding the costs of the up-interpolation layers, attention layers, and convolution layers, while the second step comes from $n\gg d$. ∎

### v2

@@ Appendix C Error Analysis of VQVAE Decoder
Lemma C.1 (Error analysis of Convolution Layer).
@@
  • Let $R>1$.
  • Assume the value of each entry in matrices can be bounded by $R$.
@@
$\|\phi_{\rm conv}(X)-\phi_{\rm conv}(X^{\prime})\|_{\infty}\leq C\epsilon R$
@@
$\leq$ $~{}\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}\epsilon\cdot R$
$\leq$ $~{}9\cdot c_{\rm in}\epsilon R$
$=$ $~{}C\epsilon R$
@@
Lemma C.2 (Error analysis of Fast VQ-VAE Decoder ).
@@
By Lemma B.6, we have shown that fast attention computation will introduce an approximation error no more than $1/\operatorname{poly}(n)$.
Then, by Lemma B.5 and Lemma C.1, we still can bounded have shown the approximation error by $1/\operatorname{poly}(n)$ after passing another up-interpolation layer, convolutions layer.
@@
Appendix D Running Time
@@
In this section, we conduct the running time analysis of every component of the $\operatorname{\mathsf{VAR}}$ model and the fast $\operatorname{\mathsf{VAR}}$ model. In Section D.1, we conduct the running time analysis of the $\operatorname{\mathsf{VAR}}$ transformer and the fast transformer. In Section D.2, we conduct the running time analysis of the feature map reconstruction layer. In Section D.3, we conduct the running time analysis of the VQVAE Decoder and fast VQVAE Decoder.
D.1 Phase 1: Running Time of Token Maps Generation
In this section, we present lemmas on the time complexity of $\operatorname{\mathsf{VAR}}$ transformer defined in Definition 3.5 and fast $\operatorname{\mathsf{VAR}}$ transformer defined in Definition 4.3.
Lemma D.1 (Running time of $\operatorname{\mathsf{VAR}}$ Transformer).
If the following conditions hold:
  • Assume the $\operatorname{\mathsf{VAR}}$ transformer defined in Definition 3.5 has $K$ attention layers.
  • Let $k_{1}\in[K]$ and $k_{2}\in[K-1]$.
  • Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k_{1}\in[K]$, the $k_{1}$-th token map $r_{k_{1}}\in\mathbb{R}^{\alpha^{k_{1}-1}\times\alpha^{k_{1}-1}\times d}$.
  • Let $\Phi_{\mathrm{up},k_{2}}$ denote the $k_{2}$-th pyramid up-interpolation layer defined in Def 3.3.
  • Let $\mathsf{Attn}_{k_{1}}$ denote the $k_{1}$-th attention layer defined in Definition 3.4.
  • Let $d=O(\log(n))$ denote the embedding size of each token.
then the time complexity of $\operatorname{\mathsf{VAR}}$ transformer $\mathsf{TF}$ is $O(n^{4+o(1)})$.
Proof.
The runtime computation of the $\operatorname{\mathsf{VAR}}$ transformer can be divided into two components: the runtime of the Attention layers $\mathsf{Attn}_{1},\dots,\mathsf{Attn}_{K}$ and the runtime of the pyramid up-interpolation layer $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$.
Part 1. Running time of the attention layers $\mathsf{Attn}_{1},\dots,\mathsf{Attn}_{K}$.
Firstly, we consider the $k_{1}$-th autoregressive token map generation. We use $L_{k_{1}}$ to denote the total number of the tokens input into the $\operatorname{\mathsf{VAR}}$ attention layer at $k_{1}$-th step and $L_{k_{1}}$ can be calculated as the following:
$L_{k_{1}}=$ $~{}\sum_{i=1}^{k_{1}}(\alpha^{i-1})^{2}$
$=$ $~{}\frac{\alpha^{2k_{1}}-1}{\alpha^{2}-1}$
$\leq$ $~{}\frac{\alpha^{2k_{1}}}{\alpha^{2}-1}$
$\leq$ $~{}\frac{\alpha^{2k_{1}}}{0.5\alpha^{2}}$
$=$ $~{}2\cdot\alpha^{2k_{1}-2}$ (5)
Thus, at the $k_{1}$-th step, we use $X_{k_{1}}\in\mathbb{R}^{L_{k_{1}}\times d}$ to denote the input matrix. The attention computation cost at the $k_{1}$-th step is $O(L_{k_{1}}^{2}d)$. We then sum up the computation time across all $K$ steps:
${\cal T}_{\mathrm{attn}}=$ $~{}O((L_{1}^{2}+L_{2}^{2}+\dots+L_{K}^{2})\cdot d)$
$\leq$ $~{}O(\sum_{k_{1}=1}^{(\log_{\alpha}{n})+1}(2\cdot\alpha^{2k_{1}-2})^{2}\cdot d)$
$=$ $~{}O(\sum_{k_{1}=1}^{(\log_{\alpha}{n})+1}4\alpha^{4k_{1}-4}\cdot d)$
$=$ $~{}O(n^{4}d)$
$=$ $~{}O(n^{4+o(1)})$
In the first step, the total time is calculated by summing the attention computation times for each $k_{1}$-th step. The second step follows directly from Eq. (D.1), while the third and fourth step is a result of basic algebraic manipulation. The last step follows from the condition that $d=O(\log(n))$.
Part 2. Running time of the pyramid up-interpolation layers $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$.
We begin by considering the runtime of the last pyramid up-interpolation layer, $\Phi_{\mathrm{up},K-1}$. From Definition 3.5, we know that the output of $\Phi_{\mathrm{up},K-1}$ serves as the input to $\mathsf{Attn}_{K}$. Therefore, the number of tokens generated by $\Phi_{\mathrm{up},K-1}$ is $L_{K}\leq 2n^{2}$, with the inequality following from Eq. D.1. Furthermore, by Eq. (1), we know that every token generated by $\Phi_{\mathrm{up},K-1}$ needs $O(d)$ times multiplications and $O(d)$ times additions. Thus, the running time of $\Phi_{\mathrm{up},K-1}$ is
$\mathcal{T}^{K-1}_{\mathrm{up}}\leq$ $~{}O(d)\cdot L_{K}$
$=$ $~{}O(n^{2}d)$ (6)
where the step comes from summing up the time cost for each token generated by $\Phi_{\mathrm{up},K-1}$.
For each $k^{\prime}\in[K-2]$, the number of tokens generated by $\Phi_{\mathrm{up},k^{\prime}}$ is less than $L_{K}$ which is due to Definition 3.3. Then we can compute the total rumming time for pyramid up-interpolation layers $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$:
$\mathcal{T}_{\mathrm{up}}\leq$ $~{}(K-1)\cdot\mathcal{T}^{K-1}_{up}$
$=$ $~{}O(\log(n))\cdot O(n^{2}d)$
$=$ $~{}O(n^{2+o(1)})$
where the step comes from summing up the running time of $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$, the second step follows from $n=\alpha^{K-1}$ and Eq (D.1) and the last step follows from $d=O(\log(n))$ and simple algebra.
Thus, by summing up the $\mathcal{T}_{\mathrm{attn}}$ and $\mathcal{T}_{\mathrm{up}}$, we know that the time complexity of $\operatorname{\mathsf{VAR}}$ transformer is $O(n^{4+o(1)})$.
Lemma D.2 (Fast Attention Computation, Theorem 1.4 of [5]).
Let $\mathsf{AAttC}$ be defined as Definition 4.2. Then we have $\mathsf{AAttC}(n,d=O(\log n),R=\Theta(\sqrt{\log n}),\delta=1/\operatorname{poly}(n))$ can be solved in time $\mathcal{T}_{\rm mat}(n,n^{o(1)},d)=n^{1+o(1)}$.
Now we can apply the result Lemma D.2 to the $\operatorname{\mathsf{VAR}}$ Transformer.
@@ D.1 Phase 1: Running Time of Token Maps Generation
Lemma D.3 (Running time of Fast $\operatorname{\mathsf{VAR}}$ Transformer, formal version of Lemma 5.1).
@@
  • Assume the fast $\operatorname{\mathsf{VAR}}$ transformer defined in Definition 4.3 has $K$ attention layers.
  • Let $k_{1}\in[K]$ and $k_{2}\in[K-1]$.
  • Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k_{1}\in[K]$, the $k_{1}$-th token map $r_{k_{1}}\in\mathbb{R}^{\alpha^{k_{1}-1}\times\alpha^{k_{1}-1}\times d}$.
  • Let $d=O(\log n)$ denote the embedding size of each token.
  • Let $\Phi_{\mathrm{up},k_{2}}$ denote the $k_{2}$-th pyramid up-interpolation layer defined in Def 3.3.
  • Let $\mathsf{AAttC}_{k_{1}}$ denote the $k_{1}$-th approximate attention layer defined in Definition 4.2.
Then, the total running time of the fast $\operatorname{\mathsf{VAR}}$ transformer $\mathsf{FTF}$ can be accelerated to $O(n^{2+o(1)})$.
Proof.
The runtime computation of the fast $\operatorname{\mathsf{VAR}}$ transformer can be divided into two components: the runtime of the approximate attention layers $\mathsf{AAttC}_{1},\dots,\mathsf{AAttC}_{K}$ and the runtime of the pyramid up-interpolation layer $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$.
Part 1. Running time of the attention layers $\mathsf{AAttC}_{1},\dots,\mathsf{AAttC}_{K}$.
To generate the $k$-th token map, let the input be $X\in\mathbb{R}^{L_{k}\times d}$. And we have
$L_{k}\leq 2\cdot\alpha^{2k-2}$ (7)
where this step is a consequence of Eq. (D.1). So the transformer computation cost at the $k$-th step can be improved from $O(L_{k}^{2}d)$ to $O(L_{k}^{1+o(1)}d)$ by using the result of Lemma D.2. Then, we sum up the computation time across all $K$ steps:
${\cal T}_{\mathrm{aattc}}=$ $~{}O((L_{1}+L_{2}+\dots+L_{K})\cdot d)$
$=$ $~{}O(n^{2+o(1)}d)$
$=$ $~{}O(n^{2+o(1)})$
where the first step follows from summing up the computation time of $\mathsf{AAttC}_{1},\dots,\mathsf{AAttC}_{K}$, the second step follows from Eq. (7), the third step follows from simple algebra and the last step follows from the condition that $d=\log(n)$.
Part 2. Running time of the pyramid up-interpolation layers $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$.
The total running time for pyramid up-interpolation layers $\Phi_{\mathrm{up},1},\dots,\Phi_{\mathrm{up},K-1}$ is the same as Lemma D.1:
$\mathcal{T}_{\mathrm{up}}=O(n^{2+o(1)})$
Thus, by summing up the $\mathcal{T}_{\mathrm{aattc}}$ and $\mathcal{T}_{\mathrm{up}}$, we know that the time complexity of $\operatorname{\mathsf{VAR}}$ transformer is $O(n^{2+o(1)})$.
D.2 Phase 2: Running Time of Feature Map Reconstruction
In this section, we analyze the total runtime of the $\operatorname{\mathsf{VAR}}$ models for feature map reconstruction.
Lemma D.4 (Running time of Feature Map Reconstruction Layer, formal version of Lemma 5.2).
If the following conditions hold:
  • Let $K\in\mathbb{N}$ denote the total number of the token maps.
@@ D.2 Phase 2: Running Time of Feature Map Reconstruction
  • Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times 1\times d}$ denote the initial token map.
@@
then the total runtime of the $\operatorname{\mathsf{VAR}}$ models for feature map reconstruction is $O(n^{2+o(1)})$.
@@
For each $k\in[K]$, $\operatorname{\mathsf{VAR}}$ Model will up-interpolate token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$ to $r^{\prime}_{k}\in\mathbb{R}^{n\times n\times d}$ by using bicubic interpolation defined in Definition 3.2. Specifically, the computation of each token in $r^{\prime}_{k}$ requires $O(d)$ multiplications and $O(d)$ additions which is due to Eq. (1). Thus, the computation cost for the up-interpolation of each token map is
${\cal T}^{k}_{\rm up}=$ $~{}O(d)\cdot n^{2}$
@@
There are total $K=O(\log n)$ token maps needed to be up-interpolated, so the total time for up-interpolation is
@@
where the first step follows from summing up the running time of $K$ scale token maps, and the second step follows from simple algebra.
Furthermore, to address the information loss in the up-interpolation process, the $\operatorname{\mathsf{VAR}}$ Model uses a convolution operation $\phi_{\mathrm{conv}}(\cdot)$ on the token map $\{r^{\prime}_{1},\dots,r^{\prime}_{K}\}$ generated by up-interpolation. We assume the convolution kernel size is $3\times 3\times d$, and the convolution layer does not change the dimension of each token map, i.e., for each $i\in[K]$, $\phi(r^{\prime}_{i})\in\mathbb{R}^{n\times n\times d}$. Hence, for every entry in $\phi(r^{\prime}_{i})$, it needs $O(d)$ operations. Then, we can have the convolution computation time for one token map is
@@
$=$ $~{}O(n^{2+o(1)})$
The first step is derived by summing the times for up-interpolation operations, convolution operations, and token map additions, while the second step is due to basic algebra. The last step follows from $d=O(\log n)$. ∎
D.3 Phase 3: Running Time of VQ-VAE Decoder
@@
Lemma D.5 (Running time of VQ-VAE Decoder).
If the following conditions hold:
@@
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th convolution layer’s output $\phi^{i}_{\rm conv}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
then the total running time of the VQ-VAE Decoder is $O(n^{4+o(1)})$.
@@
Part 1. Running time of Up-interpolation layers. For each $l\in[k_{1}]$, we assume $M^{l}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the output feature map from the $l$-th up-interpolation layer. For every token of $M^{l}$, it needs $O(d)$ multiplications and $O(d)$ additions (see more details in Definition 3.2). Thus, the computation cost for the feature map $M^{l}$ is
${\cal T}^{l}_{\rm up}=$ $~{}O(d)\cdot O(n)\cdot O(n)$
@@
$=$ $~{}O(n^{2+o(1)})$
The first step is derived by summing the computation costs for each entry of $M^{l}$, while the second step is due to basic algebra. The last step follows from the condition that $d=O(\log n)$.
@@
${\cal T}^{l}_{\rm attn}=$ $~{}O(n^{4}d)$
$=$ $~{}O(n^{4+o(1)})$
where the first step follows from computation cost for standard attention computation, the second step follows from the condition $d=O(\log n)$.
@@
${\cal T}_{\rm attn}=$ $~{}k_{2}\cdot O(n^{4+o(1)})$
$=$ $~{}O(n^{4+o(1)})$
The first step is derived by summing the computation costs for each attention layer, while the second step is due to the fact that $k_{2}$ is a constant.
@@
$=$ $~{}O(n^{2+o(1)})$
The first step is derived by summing the computation costs for each entry of $M^{l}$, while the second step stems from basic algebra. The last step follows from the condition that $d=O(\log n)$.
@@
$=$ $~{}O(n^{4+o(1)})$
The first step results from summing the computation costs of up-interpolation, attention, and convolution layers, while the second step follows from simple algebra.
@@
Lemma D.6 (Running time of Fast VQ-VAE Decoder, formal version of Lemma 5.3).
If the following conditions hold:
  • Let $k_{1},k_{2},k_{3}\in\mathbb{N}$ be constant numbers.
  • Given $X\in\mathbb{R}^{n\times n\times d}$ as the input feature map.
  • Let $d=O(\log n)$
@@
  • We assume there are $k_{2}$ approximate attention layers $\mathsf{AAttC}$ defined in Definition 4.2.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, the $i$-th approximate attention layer’s output $\mathsf{AAttC}(M)\in\mathbb{R}^{h\times w\times d}$.
@@
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th convolution layer’s output $\phi^{i}_{\rm conv}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
then the total runtime of the VQ-VAE decoder can be accelerated to $O(n^{2+o(1)})$.
@@
As the same in Eq. (D.5) and Eq. (D.5), the computation cost for up-interpolation layers and convolution layers in VQ-VAE decoder still needs $O(n^{2+o(1)}d)$.
For each $l\in[k_{2}]$, we assume $M^{l-1}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the input feature map for the $l$-th approximate attention layer. We can consider the input of the attention layer as a sequence with length $O(n^{2})$ and embedding dimension $O(d)$. By using the result of Lemma D.2, the computation cost of $M^{l}$ can be speed up to
${\cal T}^{l}_{\rm attn}=$ $~{}O(n^{2+o(1)}d)$
$=$ $~{}O(n^{2+o(1)})$
where the second step follows from the condition that $d=O(\log n)$.
@@
$=$ $~{}O(n^{2+o(1)})$
The computation cost in the first step is obtained by adding the costs of the up-interpolation layers, attention layers, and convolution layers, while the second step comes from simple algebra. ∎